I need help with this ....please!!!!
(5^-1+x^-1)/(x^2-25^-1)

Found 2 solutions by Nate, Prithwis:
Answer by Nate(3500)

(5^-1+x^-1)/(x^2-25^-1)
(5+x)^-1/(x^2-25^-1)
1/((5+x)(x^2-(1/25)))
1/(x^3-(x/25)+5x^2-(1/5))

Answer by Prithwis(166)

(5^-1+x^-1)/(x^2-25^-1)
= (1/5 + 1/x)/[(x^2 - (1/5)^2]
= [(x+5)/5x]/[(x+1/5)(x-1/5)]
= (x+5)/[(5x){(5x+1)/5}{(5x-1)/5}]
= 25(x+5)/5x(5x+1)(5x-1)
= 5(x+5)/x(x+1)(x-1)
The question may have been copied incorrectly from the book. The question makes better sense if "x^2" in the denominator is wanted as "x^-2".